a man driving his car at a constant speed of 42mph suddenly sees a cow crossing the road 60 feet ahead. at what constant deceleration (in feet per sec) is required to avoid hitting the animal. assume the reaction time of 0.3 sec before the man applies the brake.

Grade:12th pass

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please convert all the numerals in same units preferably SI

v^2 = u^2 + 2aS

v=0
u, S are given
a can be found

now
v = u+at

t can be found

Pleasedo not forget to take reaction time delay into consideration ie distance travelled at 42 mph for 0.3 sec

Regadrs